# Models That Prove Their Own Correctness

### v1

@@ Abstract
Self-Proving models satisfy that, with high probability over a random input, the model generates a correct output and successfully proves its correctness to $V\!$. The soundness property of $V$ guarantees that, for every input, no model can convince $V$ of the correctness of an incorrect output. Thus, a Self-Proving model proves correctness of most of its outputs, while all incorrect outputs (of any model) are detected by $V$. We devise a generic method for learning Self-Proving models, and we prove convergence bounds under certain assumptions.
@@ 1 Introduction
This paper initiates the study of Self-Proving models (Fig. 1) that prove the correctness of their answers via an Interactive Proof system (Goldwasser et al., 1985). Self-Proving models successfully convince a verification algorithm $V$ with worst-case soundness guarantees: for any question, with high probability over the interaction, $V$ will not be convinced of an incorrect answer. This is even when the prover with which $V$ is interacting has access to $V$’s specification, and far more computational power.
@@
  • We empirically study the efficacy of TL and Annotated-TL (ATL) for training Self-Proving transformers that compute the Greatest Common Divisor (GCD) of two integers. Table 1 demonstrates the efficacy of our methods, with additional experiments in Section 5. Our results may be of independent interest for research on the arithmetic capabilities of transformers (e.g. Charton (2024); Lee et al. (2024)). Code, data and models will be made available upon publication.
@@
Table 1: Self-Proving transformers computing the GCD. We train a 6.7M parameter GPT to compute the GCD of two integers sampled log-uniformly from $[10^{4}]$. Vanilla GPT correctly generates the GCD for almost all inputs, but does not prove correctness to a simple verification algorithm. GPT trained with Transcript Learning (GPT+TL) proves its answer 60.3% of the time; training with Annotated Transcript Learning (GPT+ATL) increases this to 96.7%. See Section 5 for more details.
@@ 2 Related Work
We conduct ablation experiments to find two deciding factors in ATL. First, we study the effect of the amount of annotation given in the form of intermediate steps (Lee et al., 2024), which is related to (AR) length complexity (Malach, 2023). Second, we characterize ATL efficacy in terms of an algebraic property of the tokenization scheme (cf. Nogueira et al. 2021; Charton 2022, 2024).
@@ 3.1 Preliminaries
Let $\Sigma$ be a set of finite tokens and $\Sigma^{\ast}$ denote the set of finite sequences of such tokens. We consider sequence-to-sequence models $F_{\theta}\colon\Sigma^{\ast}\to\Sigma^{\ast}$, which are total functions that produce an output for each possible sequence. A model is parameterized by a real-valued, finite dimensional vector $\theta$. We consider models as randomized functions, meaning that $F_{\theta}(x)$ is a random variable over $\Sigma^{\ast}$, of which samples are denoted $y\sim F_{\theta}(x)$.
@@
Let $\mu$ be a distribution of input sequences in $\Sigma^{\ast}$ and let $F^{\ast}\colon\Sigma^{\ast}\to\Sigma^{\ast}$ be a fixed (deterministic) ground-truth function. The correctness of model $F_{\theta}$ with respect to $\mu$ is
$\mathrm{corr}(\theta)\coloneqq\mathrm{corr}(F_{\theta})\coloneqq\Pr_{\begin{subarray}{c}x\sim\mu\\ y\sim F_{\theta}(x)\end{subarray}}[y=F^{\ast}(x)].$
We say that model $F_{\theta}$ is $\alpha$-correct (on average) if $\mathrm{corr}(\theta)\geq\alpha$.
@@
Given an input $x\in\Sigma^{\ast}$, the model $F_{\theta}$ “claims” that $y\sim F_{\theta}(x)$ is correct. We now define what it means to prove this claim. We will use $P_{\theta}$ to denote Self-Proving models, noting that they are formally the same object (a randomized mapping from $\Sigma^{\ast}$ to $\Sigma^{\ast}$) as vanilla models. We change the notation to emphasize that $P_{\theta}$ outputs $y\sim P_{\theta}(x)$ but can also be prompted by the verifier, unlike $F_{\theta}$ who is only expected to generate an output.
A Self-Proving model proves that $y\sim P_{\theta}(x)$ is correct to a verifier $V$ over the course of $R$ rounds of interaction (Figure 1). In each round $i\in[R]$, verifier $V$ queries $P_{\theta}$ on a sequence $q_{i}\in\Sigma^{\ast}$ to obtain an answer $a_{i}\in\Sigma^{\ast}$; once the interaction is over, $V$ accepts or rejects. For fixed $x,y\in\Sigma^{\ast}$, the decision of $V$ after interacting with $P_{\theta}$ is a random variable over $V$’s decision (accept/reject), determined by the randomness of $V$ and $P_{\theta}$. The decision random variable is denoted by $V^{P_{\theta}}(x,y)$.
@@
Fix a soundness error $s\in(0,1)$, a finite set of tokens $\Sigma$ and a ground truth $F^{\ast}\colon\Sigma^{\ast}\to\Sigma^{\ast}$. A verifier $V$ (in an Interactive Proof) for $F^{\ast}$ is a probabilistic polynomial-time algorithm that is given explicit inputs $x,y\in\Sigma^{\ast}$ and black-box (oracle) query access to a prover $P$ [^1] We intentionally write $P$ rather than $P_{\theta}$: Interactive Proofs are defined with respect to all possible provers, not just parameterized ones. It interacts with $P$ over $R$ rounds (see Figure 1) and outputs a decision $V^{P}(x,y)\in\{0,1\}$. Verifier $V$ satisfies the following two guarantees:
@@
    $\Pr[V^{P^{\ast}}\!(x,F^{\ast}(x))\ \mathrm{accepts}]=1,$
@@
    $\Pr[V^{P}(x,y)\ \mathrm{accepts}]\leq s.$
@@
We remark that, as depicted in Figure 1, each of the model’s answers depends on all previous queries and answers in the interaction. This captures the setting stateful models, e.g. a session with a chatbot.
In anticipation of Self-Proving models (Section 3.2), let us observe the following. Completeness and soundness are worst-case guarantees, meaning that they hold for all possible inputs $x\in\Sigma^{\ast}$. In particular, completeness implies that for all $x\in\Sigma^{\ast}$, the honest prover $P^{\ast}$ convinces $V$ of the correctness of $F^{\ast}(x)$; in classical proof systems there is no guarantee that an “almost honest” prover can convince the verifier (cf. Paradise (2021)). Yet, if we are to learn a prover $P_{\theta}$, we cannot expect it to agree with $P^{\ast}$ perfectly, nor can we expect it to always output $F^{\ast}(x)$. Indeed, Self-Proving models will have a distributional guarantee with respect to inputs $x\sim\mu$.
@@ 3.2 Self-Proving models
$\mathrm{ver}_{V,\mu}(\theta)\coloneqq\Pr_{\begin{subarray}{c}x\sim\mu\\ y\sim P_{\theta}(x)\end{subarray}}\left[V^{P_{\theta}}(x,y)\ \mathrm{accepts}\right].$ (1)
@@
Now, consider any input distribution $\mu$, ground-truth $F^{\ast}$, and a verifier $V$ for $F^{\ast}$ with soundness error $s$. By a union bound, if model $P_{\theta}$ is $\beta$-Verifiable, then it is $(\beta-s)$-correct. That is to say, Verifiability is formally a stronger guarantee than correctness when $V$ has small soundness error $s$.
The benefit of Verifiability over correctness is captured by the following scenario. Alice wishes to use a model $P_{\theta}$ to compute some functionality $F^{\ast}$ on an input $x_{0}$ in a high risk setting. Alice generates $y_{0}\sim P_{\theta}(x_{0})$. Should Alice trust that $y_{0}$ is correct? If Alice has a held-out set of labeled samples she can estimate $P_{\theta}$’s average correctness on $\mu$. Unfortunately, (average) correctness provides no guarantee regarding for the correctness of the particular $(x_{0},y_{0})$ that Alice has in hand. If, however, Alice has access to a verifier $V$ for which $P_{\theta}$ is Self-Proving, then she can trust the model on an input-by-input (rather than average-case) basis: Alice can execute $V$ on $(x_{0},y_{0})$ and black-box access to $P_{\theta}$. Soundness of $V$ guarantees that if $y_{0}$ is incorrect, then $V$ rejects with high probability, in which case Alice should either generate $P_{\theta}(x_{0})$ again—or find a better model.
@@ 4 Learning Self-Proving Autoregressive Models
The challenge is to align the model with a verifier. We assume that the learner has access to input samples $x\sim\mu$ and correct outputs $F^{\ast}(x)$, as well as the verifier specification (code). Additionally, the learner can emulate the verifier, as the latter is required to be computationally efficient. [^3] We refer the reader to classical literature on Interactive Proof systems for formal definitions of computational efficiency (e.g. Goldreich (2008)).
@@
Our general approach is inspired by Reinforcement Learning from Human Feedback (Christiano et al., 2017), a method for aligning models with human preferences, which has recently been used to align sequence-to-sequence models (Ouyang et al., 2022). However, there are two important differences between humans and algorithmic verifiers: (1) Verifiers are efficient algorithms which may be emulated by the learner. This is unlike humans, whose preferences are costly to obtain. On the other hand, (2) verifiers make a single-bit decision at the end of an interaction, but cannot guide the prover (model) in intermediate rounds. In RL terms, this is known as the exploration problem for sparse reward signals (e.g., Ladosz et al. 2022).
@@ 4.1 Transcript Learning
We present an algorithm for learning Self-Proving models which uses access to a distribution of accepting transcripts. This is a reasonable assumption to make when the honest prover $P^{\ast}$ (see Definition 3.2) is efficient, as is the case in Doubly-Efficient Interactive Proof systems as defined by Goldwasser et al. (2015) and developed in other theoretical (e.g. Goldreich and Rothblum 2018) and applied (e.g. Zhang et al. 2021) works. In this case, an honest prover $P^{\ast}$ can be run by the learner during training to collect accepting transcripts without incurring heavy computational cost.
@@
Transcript Learning (TL) trains a Self-Provable model by autoregressively optimizing towards generating accepting transcripts. It is described in Algorithm 1. At a very high level, it works by repeatedly sampling $x\sim\mu$ and transcript $y^{\ast}\pi^{\ast}\sim\mathcal{T}^{\ast}(x)$, and updating the logits $\log p_{\theta}$ towards agreeing with $y^{\ast}\pi^{\ast}$ via Gradient Ascent. We prove that, under certain conditions, it is expected to output a Self-Provable model. We fully describe the conditions in Section A.1; we provide an informal statement (with an undefined “expected agreement” quantity) next.
@@
Fix an input distribution $\mu$, a verifier $V$ and an autoregressive model family $\{P_{\theta}\}_{\theta}$. Fix a transcript generator $\mathcal{T}^{\ast}$ such that expected agreement $\mathcal{T}^{\ast}$ is convex in $\theta$. For any $\varepsilon>0$ such that there exists $\theta^{\ast}$ with at least $\geq 1-\varepsilon/2$ expected agreement with $\mathcal{T}^{\ast}$, let $B$ be the minimal norm of such $\theta^{\ast}$. Let $\rho>0$ such that for all $\theta$ with $||\theta||<B$, the logits $\log p_{\theta}$ are $\rho$-Lipschitz in $\theta$. Denote by $\bar{\theta}$ the output of TL running for number of iterations
$N=R^{2}\cdot(L_{a}+1)^{2}\cdot\frac{4\rho^{2}\cdot B^{2}}{\epsilon^{2}}$ (2)
and learning rate $\lambda=\varepsilon/2R^{2}L_{a}^{2}\rho$. Then the expected Verifiability of $\bar{\theta}$ is at least $1-\varepsilon$.
The proof works by reduction to Stochastic Gradient Descent (SGD). The main challenge is to prove that the learner can use its only available tools—sampling honest transcripts, emulating the verifier, and differentiating the logits—to estimate the gradient of (a lower bound on) the Verifiability of $P_{\theta}$.
Looking at Equation 2, we see that the sample complexity of TL grows like that of SGD (e.g. Shalev-Shwartz and Ben-David 2014), multiplied by the number of rounds and length of answers in the proof system. Minimizing these quantities (known collectively as the communication complexity) has been an overarching goal in the study of proof systems (e.g. Goldreich and Håstad (1998); Goldreich et al. (2002); Reingold et al. (2021). Theorem 4.1 formally shows the benefit of communication-efficient proof systems in the context of Self-Proving models.

### v2

@@ Abstract
Self-Proving models satisfy that, with high probability over an input sampled from a given distribution, the model generates a correct output and successfully proves its correctness to $V\!$. The soundness property of $V$ guarantees that, for every input, no model can convince $V$ of the correctness of an incorrect output. Thus, a Self-Proving model proves correctness of most of its outputs, while all incorrect outputs (of any model) are detected by $V$. We devise a generic methods for learning Self-Proving models, and prove its convergence under certain assumptions.
@@ 1 Introduction
This paper initiates the study of Self-Proving models (Fig. 1) that prove the correctness of their answers via an Interactive Proof system (Goldwasser et al., 1985). Self-Proving models successfully convince a verification algorithm $V$ with worst-case soundness guarantees: for any question, $V$ rejects all incorrect answers with high probability over the interaction. This guarantee holds even against provers that have access to $V$’s specification, and unbounded computational power.
@@
Guarantee Type Def.
$V$ Completeness & Soundness Worst-case $\forall x,y$ 3.2
$P_{\theta}$ Verifiability Average-case $x\sim\mu$, $y\sim P_{\theta}(x)$ 3.4
Table 1: Formal guarantees. Completeness and soundness are fundamental guarantees of a verification algorithm $V$. Verifiability (novel in this work) is a feature of a model $P_{\theta}$ with respect to a verifier $V$ and input distribution $\mu$. Importantly, $V$’s soundness holds for any input $x$ and output $y$.
@@
  • We empirically study TL and Annotated-TL (ATL) for training Self-Proving transformers that compute the Greatest Common Divisor (GCD) of two integers. Table 2 demonstrates the efficacy of our methods, with additional experiments in Section 5. Our results may be of independent interest for research on the arithmetic capabilities of transformers (e.g. Charton 2024; Lee et al. 2024). Code, data and models will be made available upon publication.
Scope.
This paper contains a theory of learned models that prove their own correctness via an Interactive Proof system. The fascinating and well-studied question of which settings are verifiable in an Interactive Proof system is beyond our scope. Our theory is general in that it pertains to any such setting, e.g., any decision problem solvable in polynomial space (Shamir, 1992). See Goldreich (2008) for a primer on Proof systems more broadly.
@@
Table 2: Self-Proving transformers computing the GCD. We train a 6.7M parameter GPT to compute the GCD of two integers sampled log-uniformly from $[10^{4}]$. Vanilla GPT correctly generates the GCD for almost all inputs, but does not prove correctness to a simple verification algorithm. GPT trained with Transcript Learning (GPT+TL) proves its answer 60.3% of the time; training with Annotated Transcript Learning (GPT+ATL) increases this to 96.7%. See Section 5 for details.
@@ 2 Related Work
We conduct ablation experiments to find two deciding factors in ATL. First, we study the effect of the amount of annotation given in the form of intermediate steps (Lee et al., 2024), which is related to autoregressive length complexity (Malach, 2023). Second, we characterize ATL efficacy in terms of an algebraic property of the tokenization scheme (cf. Nogueira et al. 2021; Charton 2022, 2024).
@@ 3.1 Preliminaries
Let $\Sigma$ be a set of finite tokens and $\Sigma^{\ast}$ denote the set of finite sequences of such tokens. We consider sequence-to-sequence models $F_{\theta}\colon\Sigma^{\ast}\to\Sigma^{\ast}$, which are total functions that produce an output for each possible sequence. A model is parameterized by a real-valued, finite dimensional vector $\theta$. We consider models as randomized functions, meaning that $F_{\theta}(x)$ is a random variable over $\Sigma^{\ast}$, of which samples are denoted by $y\sim F_{\theta}(x)$.
@@
Let $\mu$ be a distribution of input sequences in $\Sigma^{\ast}$ and let $F^{\ast}\colon\Sigma^{\ast}\to\Sigma^{\ast}$ be a fixed (deterministic) ground-truth function. For any $\alpha\in[0,1]$, we say that model $F_{\theta}$ is $\alpha$-correct (with respect to $\mu$) if
$\Pr_{\begin{subarray}{c}x\sim\mu\\ y\sim F_{\theta}(x)\end{subarray}}[y=F^{\ast}(x)]\geq\alpha.$
@@
Given an input $x\in\Sigma^{\ast}$, the model $F_{\theta}$ “claims” that $y\sim F_{\theta}(x)$ is correct. We now define what it means to prove this claim. We will use $P_{\theta}$ to denote Self-Proving models, noting that they are formally the same object [^1] Both are randomized mappings from $\Sigma^{\ast}$ to $\Sigma^{\ast}$. as non-Self-Proving (“vanilla”) models $F_{\theta}$. This notational change is to emphasize that $P_{\theta}$ first outputs $y\sim P_{\theta}(x)$ and is then prompted by the verifier, unlike $F_{\theta}$ who only generates an output $y\sim F_{\theta}(x)$.
A Self-Proving model proves that $y\sim P_{\theta}(x)$ is correct to a verifier $V$ over the course of $R$ rounds of interaction (Figure 1). In each round $i\in[R]$, verifier $V$ queries $P_{\theta}$ on a sequence $q_{i}\in\Sigma^{\ast}$ to obtain an answer $a_{i}\in\Sigma^{\ast}$; once the interaction is over, $V$ accepts or rejects. For fixed $x,y\in\Sigma^{\ast}$, the decision of $V$ after interacting with $P_{\theta}$ is a random variable over $V$’s decision (accept/reject), determined by the randomness of $V$ and $P_{\theta}$. The decision random variable is denoted by $V,P_{\theta}\right>(x,y)$.
@@
Fix a soundness error $s\in(0,1)$, a finite set of tokens $\Sigma$ and a ground truth $F^{\ast}\colon\Sigma^{\ast}\to\Sigma^{\ast}$. A verifier $V$ (in an Interactive Proof) for $F^{\ast}$ is a probabilistic polynomial-time algorithm that is given explicit inputs $x,y\in\Sigma^{\ast}$ and black-box (oracle) query access to a prover $P$ [^2] We intentionally write $P$ rather than $P_{\theta}$: Interactive Proofs are defined with respect to all possible provers, not just parameterized ones. It interacts with $P$ over $R$ rounds (see Figure 1) and outputs a decision $V,P\right>(x,y)\in\{0,1\}$. Verifier $V$ satisfies the following two guarantees:
@@
    $\Pr[\left<V,P^{\ast}\right>\!(x,F^{\ast}(x))\ \mathrm{accepts}]=1,$
@@
    $\Pr[\left<V,P\right>(x,y)\ \mathrm{accepts}]\leq s,$
@@
Remark 3.3 (Verifier efficiency).
Definition 3.2 requires that $V$ is a polynomial-time algorithm whereas provers are unbounded. This captures a requirement for efficient verification. We chose polynomial time as a measure of efficiency because it is common Proof systems literature. That said, one could adapt Definition 3.2 to fit alternative efficiency measures, such as space complexity (Condon and Lipton, 1989) or circuit depth (Goldwasser et al., 2007). Regardless of which measure is taken, to avoid a trivial definition it is crucial that $V$ should be more efficient than the honest prover $P^{\ast}$; else, $V$ can simply execute $P^{\ast}$ to perform the computation itself.
@@
As depicted in Figure 1, each of the model’s answers depends on all previous queries and answers in the interaction. This captures the setting stateful models, e.g. a session with a chatbot.
Towards defining Self-Proving models (Section 3.2), let us observe the following. Completeness and soundness are worst-case guarantees, meaning that they hold for all possible inputs $x\in\Sigma^{\ast}$. In particular, completeness implies that for all $x\in\Sigma^{\ast}$, the honest prover $P^{\ast}$ convinces $V$ of the correctness of $F^{\ast}(x)$; in classical proof systems there is no guarantee that an “almost honest” prover can convince the verifier (cf. Paradise (2021)). Yet, if we are to learn a prover $P_{\theta}$, we cannot expect it to agree with $P^{\ast}$ perfectly, nor can we expect it to always output $F^{\ast}(x)$. Indeed, Self-Proving models will have a distributional guarantee with respect to inputs $x\sim\mu$.
@@ 3.2 Self-Proving models
$\mathrm{ver}_{V,\mu}(\theta)\coloneqq\Pr_{\begin{subarray}{c}x\sim\mu\\ y\sim P_{\theta}(x)\end{subarray}}\left[\left<V,P_{\theta}\right>(x,y)\ \mathrm{accepts}\right].$ (1)
@@
Remark 3.5 (Verifiability $\implies$ correctness).
Notice that the ground-truth $F^{\ast}$ does not appear in Definition 3.4 except for the first sentence. Indeed, once it is established that $V$ is a verifier for $F^{\ast}$ (as per Definition 3.2), then Verifiability w.r.t $V$ implies correctness w.r.t $F^{\ast}$: Consider any input distribution $\mu$, ground-truth $F^{\ast}$, and a verifier $V$ for $F^{\ast}$ with soundness error $s$. By a union bound, if model $P_{\theta}$ is $\beta$-Verifiable, then it is $(\beta-s)$-correct. That is to say, Verifiability is formally a stronger guarantee than correctness when $V$ has small soundness error $s$.
As depicted in Figure 1, a Self-Proving model $P_{\theta}$ plays a dual role: first, it generates an output $y\sim P_{\theta}(x)$, and then it proves the correctness of this output to $V$. Note also that Self-Provability is a feature of a model, unlike completeness and soundness which are features of a verifier (see Table 1).
The benefit of Verifiability over correctness is captured by the following scenario. Alice wishes to use a model $P_{\theta}$ to compute some functionality $F^{\ast}$ on an input $x_{0}$ in a high risk setting. Alice generates $y_{0}\sim P_{\theta}(x_{0})$. Should Alice trust that $y_{0}$ is correct? If Alice has a held-out set of labeled samples, she can estimate $P_{\theta}$’s average correctness on $\mu$. Unfortunately, (average) correctness provides no guarantee regarding for the correctness of the particular $(x_{0},y_{0})$ that Alice has in hand. If, however, Alice has access to a verifier $V$ for which $P_{\theta}$ is Self-Proving, then she can trust the model on an input-by-input (rather than average-case) basis: Alice can execute $V$ on $(x_{0},y_{0})$ and black-box access to $P_{\theta}$. Soundness of $V$ guarantees that if $y_{0}$ is incorrect, then $V$ rejects with high probability, in which case Alice should either generate $P_{\theta}(x_{0})$ again—or find a better model.
@@ 4 Learning Self-Proving Autoregressive Models
The challenge is to align the model with a verifier. We assume that the learner has access to input samples $x\sim\mu$ and correct outputs $F^{\ast}(x)$, as well as the verifier specification (code). Additionally, the learner can emulate the verifier, as the latter is computationally efficient (Remark 3.3).
@@
Our general approach is inspired by Reinforcement Learning from Human Feedback (Christiano et al., 2017), a method for aligning models with human preferences, which has recently been used to align sequence-to-sequence models (Ouyang et al., 2022). However, there are two important differences between humans and algorithmic verifiers: (1) Verifiers are efficient algorithms which may be emulated by the learner. This is unlike humans, whose preferences are costly to obtain. On the other hand, (2) verifiers make a single-bit decision at the end of an interaction, but cannot guide the prover (model) in intermediate rounds. In RL terms, this is known as the exploration problem for sparse reward signals (e.g. Ladosz et al. 2022).
@@ 4.1 Transcript Learning
We present an algorithm for learning Self-Proving models which uses access to a distribution of accepting transcripts. This is a reasonable assumption to make when the honest prover $P^{\ast}$ (see Definition 3.2) is efficient, as in the case in Doubly-Efficient Interactive Proof systems as defined by Goldwasser et al. (2015) and developed in other theoretical (e.g. Goldreich and Rothblum 2018) and applied (e.g. Zhang et al. 2021) works. In this case, an honest prover $P^{\ast}$ can be run by the learner during training to collect accepting transcripts without incurring heavy computational cost.
@@
Transcript Learning (TL) trains a Self-Provable model by autoregressively optimizing towards generating accepting transcripts. It is described in Algorithm 1. At a very high level, it works by repeatedly sampling $x\sim\mu$ and transcript $y^{\ast}\pi^{\ast}\sim\mathcal{T}^{\ast}(x)$, and updating the logits $\log p_{\theta}$ towards agreeing with $y^{\ast}\pi^{\ast}$ via Gradient Ascent. We prove that, under certain conditions, it is expected to output a Self-Provable model.
@@
Fix an input distribution $\mu$, a verifier $V$, a transcript generator $\mathcal{T}^{\ast}$, an autoregressive model family $\{P_{\theta}\}_{\theta}$ parameterized by $\theta\in\mathbb{R}^{d}$ for some $d\in\mathbb{N}$, and a norm |||||$\cdot$||| | ⋅ | | on $\mathbb{R}^{d}$. Assume that the agreement function $A\colon\mathbb{R}^{d}\to[0,1]$ defined by
$A(\theta)\coloneqq\Pr_{\begin{subarray}{c}x\sim\mu\\ {\pi^{\ast}\sim\mathcal{T}^{\ast}(x)}\end{subarray}}[\mathrm{Transcript}(\left<V,P_{\theta}\right>(x))=\pi^{\ast}]$
is concave and differentiable in $\theta$. For any $\varepsilon>0$, let $B_{\mathrm{Norm}}$, $B_{\mathrm{Lip}}$ and $C$ be upper-bounds such that the following conditions hold.
  • There exists $\theta^{\ast}\in\mathbb{R}^{d}$ with $||\theta^{\ast}||<B_{\mathrm{Norm}}$ such that $A(\theta^{\ast})\geq 1-\varepsilon/2$.
  • For all $\theta$, the logits of $P_{\theta}$ are $B_{\mathrm{Lip}}$-Lipschitz in $\theta$.
  • The total number of tokens sent by the prover to the verifier $V$ in any interaction is at most $C$.
Denote by $\bar{\theta}$ the output of TL running for number of iterations
$N\geq 4\cdot C^{2}\cdot\frac{B_{\mathrm{Norm}}^{2}\cdot B_{\mathrm{Lip}}^{2}}{\varepsilon^{2}}$ (2)
and learning rate $\lambda=B_{\mathrm{Norm}}/CB_{\mathrm{Lip}}\sqrt{N}$. Then the expected Verifiability of $\bar{\theta}$ is at least $1-\varepsilon$.
The proof (Appendix A) goes by reduction to Stochastic Gradient Descent (SGD). We show that the learner can use only its available tools—sampling honest transcripts, emulating the verifier, and differentiating the logits—to estimate the agreement gradient $\nabla A(\theta)$. Since the agreement $A(\theta)$ lower bounds the Verifiability of $P_{\theta}$, the former can be used as a surrogate objective for the latter.
The conditions for Theorem 4.1 can be split into two. First, the standard conditions used to prove SGD convergence: convexity, [^5] Theorem 4.1 requires concavity because it guarantees maximization, rather than minimization. We leave it for future work to relax the differentiability assumption. $B_{\mathrm{Norm}}$-boundedness, and $B_{\mathrm{Lip}}$-Lipschitzness. Second, there is a bound $C$ on the communication complexity of the prover in the Interactive Proof system.
Quantitatively, the efficiency of TL is captured by the number of iterations $N$. It is desirable to minimize $N$, which is also the number of samples needed from the distribution $\mu$ and the transcript generator $\mathcal{T}^{\ast}$. Like the conditions on the theorem, the bound on $N$ can too be decomposed into two factors: The right factor is the complexity of SGD ($B_{\mathrm{Norm}}^{2}B_{\mathrm{Lip}}^{2}/\varepsilon^{2}$), and the left factor $O(C^{2})$ is the communication complexity of the proof system. Minimizing communication complexity has been an overarching goal in the study of proof systems (e.g. Goldreich and Håstad 1998; Goldreich et al. 2002; Reingold et al. 2021). Theorem 4.1 formally shows the benefit of communication-efficient proof systems in the context of Self-Proving models.